OpenStudy (anonymous):

Find the length of the radius of the given circle. (x - 5)2 + (y + 3)2 = 25

OpenStudy (anonymous):

you have \((x-h)^2+(y-k)^2=r^2\) the equation for a circle with radius \(r\) since \(r^2=25\) you know \(r=5\)

OpenStudy (anonymous):

thx man hey how do you find the coordinates of the center of the given circle. do you use the same process?

OpenStudy (precal):

(5, -3) is the center

OpenStudy (anonymous):

thx

OpenStudy (precal):

anytime :)